40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 327     Entry time: Thu Feb 21 09:56:26 2008
Author: steve 
Type: Update 
Category: PSL 
Subject: the laser is back 
The laser and the psl recovered.
The water chiller temp is 19.98C and head temp 18.4C
Power 3.2W
PMC_T 3.1

c1iovme was restarted and the mc is locking now
Attachment 1: laserisback.jpg  105 kB  | Hide | Hide all
laserisback.jpg
ELOG V3.1.3-